# Learning to Call: A Field Trial of a Collaborative Bandit Algorithm for Improved Message Delivery in Mobile Maternal Health

### v1

@@ 1 Introduction
Maternal health remains a critical public health concern in India and various other developing countries globally (MomConnect (2023), Nshimirimana et al. (2012), Hategeka and Law (2019), Ward et al. (2020), mHealth Tanzania (2013)), with millions of women having limited access to timely and accurate information during pregnancy and postpartum. Mobile health (mHealth) programs (Murthy et al. (2020a), Chowdhury et al. (2019), Kabukye et al. (2021)) that use automated voice messages, have the ability to deliver such critical maternal and child health information. Recognizing the need for information and the promise of mHealth programs, the Government of India launched the Kilkari program, a nationwide mobile health initiative that delivers weekly voice messages that contain essential maternal health information to more than 10 million registered mothers  [^1] https://rchrpt.mohfw.gov.in/RCHRPT/Kilkari/Kilkari_Message.aspx. mHealth programs such as these play a vital role in reducing maternal mortality rates - a key target within the WHO’s Sustainable Development Goals WHO . These messages cover vital topics such as iron and calcium supplementation, antenatal care, and postnatal practices, aiming to improve maternal health outcomes throughout the country.
However, the effectiveness of this large-scale program is contingent upon successful message delivery. Currently, Kilkari employs a random call scheduling strategy, attempting to reach mothers, with up to nine re-attempts (until the call is picked up), but without considering individual preferences for call timing. This approach often results in missed calls, crucial bandwidth spent on re-attempts and most importantly limiting the reach and impact of crucial health information JJH et al. (2021); Mohan et al. (2021); Lalan et al. (2023). To address this challenge, recently, a stochastic bandit approach was proposed to learn appropriate timing for calls to mothers. Whereas learning individually the appropriate timing to call each mother is expensive, a collaborative bandit approach attempts to harness similarity among the mothers to jointly learn their preferences for call timings Pal et al. (2024). Whereas this approach has shown promise in simulations, its performance in real-world field trials remains unknown. To address this limitation, this paper presents a field trial of the collaborative bandit algorithm Pal et al. (2024) designed to optimize call scheduling by learning mothers’ preferred call times.
@@
This research contributes to the growing body of literature in the application of machine learning in mobile health interventions Verma et al. (2023); Mate et al. (2022); Nair et al. (2022). By demonstrating the effectiveness of a collaborative bandit algorithm in a real-world setting, we highlight the potential for personalized call scheduling to enhance the reach and impact of maternal health programs at scale. Given the national scope of Kilkari and the potential for improved message delivery to millions of beneficiaries, our findings have significant implications for public health policy and practice in India and beyond.
@@
This work was conducted as a joint effort between a research team from a non-profit in India called ARMMAN (arm ) and Google Deepmind India a non-profit organization in India as reflected in the co-authorship of this paper. It is crucial to highlight that the beneficiary data utilized in this research is fully anonymized, and no socio-demographic features were available to the research team. To ensure data privacy and security, the experimental infrastructure was managed exclusively by the ARMMAN team, who were the only individuals with access to the raw beneficiary data. The Google Deepmind researchers contributed by advising the ARMMAN team on the collaborative bandit algorithm, specifically the algorithm’s implementation and subsequently collaborating on the analysis of the resulting study. The ARMMAN team has followed general guidelines related to ethics approvals laid down by Indian Council for Medical Research (ICMR).
@@ 2 Related Work
mHealth programs provide essential health information though automated voice messages to a large number of beneficiaries Hegde and Doshi (2016); Murthy et al. (2020b), which implies any improvement to the program positively affects a lot of beneficiaries. Previously, AI has been applied to schedule interventions Mate et al. (2022); Verma et al. (2023) and showed positive behavioral outcomes Dasgupta et al. (2024). LeFevre et al. (2019) talk about the protocol for an individually controlled randomized control trial in an attempt to show the effectiveness of Kilkari.
Scheduling the time of the day the beneficiaries are called using collaborative bandits Pal et al. (2024) showed promise in simulation and this work aims to test it out in a pilot study. The analysis in Bashingwa et al. (2021b) shows that there are preferred slots for calling beneficiaries. They further show that most calls which are picked are done so by the third attempt. These factors point towards the advantages of scheduling these calls in a non-random manner.
Multi-armed bandits represent a well-researched and potent approach for tackling diverse resource allocation challenges. Numerous methodologies, including phased elimination Lattimore and Szepesvári (2020); Slivkins et al. (2019), Upper Confidence Bound (UCB) Auer et al. (2002), Thompson Sampling Thompson (1933); Agrawal and Goyal (2012), and Best-arm Identification Agrawal et al. (2020); Garivier and Kaufmann (2016), have undergone thorough investigation. The collaborative bandit problem has witnessed a surge in interest recently, driven by the widespread adoption of recommender systems Bresler et al. (2016); Dadkhahi and Negahban (2018). Under specific conditions, several algorithms with robust theoretical guarantees have been developed Pal et al. (2023a); Jain and Pal (2022). An algorithm suited for scenarios with approximate low-rank structure, was introduced by Pal et al. (2024) and is evaluated in this field study.
@@ 3.2 Collaborative Bandits
Pal et al. (2024) address the challenge of optimizing time slot selection in mobile health programs like Kilkari, where the goal is to deliver automated voice messages to beneficiaries at times they are most likely to engage. To tackle this, Pal et al. (2024) formulate the problem as a multi-agent multi-armed bandit problem Slivkins et al. (2019). Here, each beneficiary is modeled as an agent, and each possible time slot for delivering the message is considered an arm. The key idea is to learn the preferences of these agents (beneficiaries) for different arms (time slots) through repeated interactions (i.e., attempting to deliver messages).
To efficiently solve this multi-agent bandit problem, the underlying algorithm in Pal et al. (2023b, a) is applied to the current problem. This framework leverages the assumption that the preferences of beneficiaries are not entirely independent but rather share some underlying structure. Specifically, it assumes that the matrix representing beneficiaries’ preferences (e.g., the probability of a beneficiary picking up a call at a given time slot) is approximately low-rank. This low-rank assumption implies that there are a few latent factors that explain a significant portion of the variability in beneficiaries’ time slot preferences. The collaborative bandit algorithm exploits this low-rank structure to learn more efficiently by sharing information across beneficiaries, rather than learning each beneficiary’s preferences in isolation.
Pal et al. (2024) introduces two novel algorithms: Greedy Matrix Completion (MC) and Phased MC. Phased MC is the key algorithm we use in this work. Greedy MC first has a long random phase where arms are picked randomly, followed by prediction which is followed thereafter. Phased MC operates by updating the estimates in ”phases” which implies that no lengthy exploration phase is required to obtain an initial estimate. To allow for exploration during prediction, a Boltzmann noise Cesa-Bianchi et al. (2017) is added. It also uses variance reduction techniques to improve robustness to noise. Since the exploration phase required in this algorithm is not as large, it hence prevents chances of dropout early in the trial due to random calls.
@@ 4.1 Randomization
Random Group (Control): This group received calls using the current Kilkari call retry algorithm Bashingwa et al. (2021a). This group considered $6416$ beneficiaries for the analysis.
Treatment Group (Collaborative Bandit): This group received calls scheduled using a collaborative bandit algorithm designed to learn and adapt to individual beneficiaries’ preferred call times. This group considered $6490$ beneficiaries for the analysis. The algorithm from Pal et al. (2024) relies on an offline low rank matrix completion oracle $\mathcal{O}$. For an unknown matrix $Z\in\mathbb{R}^{m\times n}$, $\mathcal{O}$ takes a subset of noisy observations of a matrix as position $\omega$ ($\{M_{ij}\}_{(i,j)\in\omega}$) as input, and returns an estimate $\hat{Z}$ of $Z$. To implement this oracle, we minimize the following nuclear norm regularized objective:
$\textrm{minimize}_{\hat{Z}}\sum_{(i,j)\in\omega}(M_{ij}-\hat{Z}_{ij})^{2}+\lambda||\hat{Z}||_{\star},$ (1)
where $\lambda>0$ is the regularization parameter and $||\cdot||_{\star}$ denotes the nuclear norm for a matrix as shown in Pal et al. (2024). The hyper-parameters of the model were tuned by holding out 20% of the most recent call data as validation, and then choosing the best hyper-parameters by calculating the score on this data based on predictions via grid search.
@@ 5 Preliminaries
In this section we mathematically define the call pick-up rates and its variants that are used for the analysis in the later section. The index $i$ represents a beneficiary (or user), $j\in[1,7]$ being one of the seven time slot IDs chosen, $t$ being the day, and $r\in[0,2]$ being the re-attempt number for that slot, i.e. $r=0$ being the first call, and $r=1$ being the second call made, if the first call wasn’t picked up. The time windows for the 7 time slot IDs are detailed in the Appendix in Table 6. Let $call$ be mapped uniquely to the tuple $(i,j,t,r)$. Let $A_{call}\equiv A_{i,j,t,r}\in\{0,1\}$ denote whether a call attempt was made for user $i$ during time-slot $j$ on day $t$ and whether it was the $r$-th re-attempt. Similarly, let $p_{call}\equiv p_{i,j,t,r}\in\{0,1\}$ denote whether an attempted call was picked or not. We assume that the set of calls $\{call|p_{call}=1\}\subseteq\{call|A_{call}=1\}$. We now define the pooled pick-up rate as,
@@ 6.6 Summary
Some of the generalizable insights learned from this work are as follows. Collaborative bandit algorithms offer a promising approach for personalized intervention delivery in mobile health By iteratively learning from user responses and interactions, these algorithms can adapt to individual preferences and maximize engagement as compared to random calling algorithms. Our goal was to evaluate the algorithm’s ability to improve call pick-up rates compared to the baseline random calling strategy. This research contributes to the growing body of literature in the application of machine learning in mobile health interventions Verma et al. (2023); Mate et al. (2022); Nair et al. (2022). By demonstrating the effectiveness of a collaborative bandit algorithm in a real-world setting, we highlight the potential for personalized call scheduling to enhance the reach and impact of maternal health programs at scale. Given the national scope of Kilkari and the potential for improved message delivery to millions of beneficiaries, our findings have significant implications for public health policy and practice in India and beyond.

### v2

# Learning to Call: A Field Trial of a Collaborative Bandit Algorithm for Optimizing Call Timing in Mobile Maternal Health

@@ 1 Introduction
Maternal health remains a critical public health concern in India and various other developing countries globally (momconnect, nshimirimana2012designing, hategeka2019effect, Ward2020-tq, mHealthTanzania2013), with millions of women having limited access to timely and accurate information during pregnancy and postpartum. Mobile health (mHealth) programs (Murthy2020, Chowdhury2019, Kabukye2021-yy) that use automated voice messages, have the ability to deliver such critical maternal and child health information. Recognizing the need for information and the promise of mHealth programs, the Government of India launched the Kilkari program, a nationwide mobile health initiative that delivers weekly voice messages that contain essential maternal health information to more than 10 million registered mothers  [^1] https://rchrpt.mohfw.gov.in/RCHRPT/Kilkari/Kilkari_Message.aspx. mHealth programs such as these play a vital role in reducing maternal mortality rates - a key target within the WHO’s Sustainable Development Goals WHOTargetMaternal. These messages cover vital topics such as iron and calcium supplementation, antenatal care, and postnatal practices, aiming to improve maternal health outcomes throughout the country.
However, the effectiveness of this large-scale program is contingent upon successful message delivery. Currently, Kilkari employs a random call scheduling strategy, attempting to reach mothers, with up to nine re-attempts (until the call is picked up), but without considering individual preferences for call timing. This approach often results in missed calls, crucial bandwidth spent on re-attempts and most importantly limiting the reach and impact of crucial health information JJH; mohan2021can; lalan2023analyzing. To address this challenge, recently, a stochastic bandit approach was proposed to learn appropriate timing for calls to mothers. Whereas learning individually the appropriate timing to call each mother is expensive, a collaborative bandit approach attempts to harness similarity among the mothers to jointly learn their preferences for call timings pal2024improving. Whereas this approach has shown promise in simulations, its performance in real-world field trials remains unknown. To address this limitation, this paper presents a field trial of the collaborative bandit algorithm pal2024improving designed to optimize call scheduling by learning mothers’ preferred call times.
@@
This research contributes to the growing body of literature in the application of machine learning in mobile health interventions verma2023deployed; mate2022field; nair2022adviser. By demonstrating the effectiveness of a collaborative bandit algorithm in a real-world setting, we highlight the potential for personalized call scheduling to enhance the reach and impact of maternal health programs at scale. Given the national scope of Kilkari and the potential for improved message delivery to millions of beneficiaries, our findings have significant implications for public health policy and practice in India and beyond.
@@
This work was conducted as a joint effort between a research team from a non-profit in India called ARMMAN (armmanArmmanHome) and Google Deepmind India a non-profit organization in India as reflected in the co-authorship of this paper. It is crucial to highlight that the beneficiary data utilized in this research is fully anonymized, and no socio-demographic features were available to the research team. To ensure data privacy and security, the experimental infrastructure was managed exclusively by the ARMMAN team, who were the only individuals with access to the raw beneficiary data. The Google Deepmind researchers contributed by advising the ARMMAN team on the collaborative bandit algorithm, specifically the algorithm’s implementation and subsequently collaborating on the analysis of the resulting study. The ARMMAN team has followed general guidelines related to ethics approvals laid down by Indian Council for Medical Research (ICMR).
@@ 2 Related Work
mHealth programs provide essential health information though automated voice messages to a large number of beneficiaries hegde2016assessing; murthy2020effects, which implies any improvement to the program positively affects a lot of beneficiaries. Previously, AI has been applied to schedule interventions mate2022field; verma2023deployed and showed positive behavioral outcomes dasgupta2024preliminary. lefevre2019stage talk about the protocol for an individually controlled randomized control trial in an attempt to show the effectiveness of Kilkari.
Scheduling the time of the day the beneficiaries are called using collaborative bandits pal2024improving showed promise in simulation and this work aims to test it out in a pilot study. The analysis in bashingwa2021assessing shows that there are preferred slots for calling beneficiaries. They further show that most calls which are picked are done so by the third attempt. These factors point towards the advantages of scheduling these calls in a non-random manner.
Multi-armed bandits represent a well-researched and potent approach for tackling diverse resource allocation challenges. Numerous methodologies, including phased elimination lattimore2020bandit; slivkins2019introduction, Upper Confidence Bound (UCB) auer2002finite, Thompson Sampling thompson1933likelihood; agrawal2012analysis, and Best-arm Identification agrawal2020optimal; garivier2016optimal, have undergone thorough investigation. The collaborative bandit problem has witnessed a surge in interest recently, driven by the widespread adoption of recommender systems bresler2016collaborative; dadkhahi2018alternating. Under specific conditions, several algorithms with robust theoretical guarantees have been developed pal2023optimal; jain2022online. An algorithm suited for scenarios with approximate low-rank structure, was introduced by pal2024improving and is evaluated in this field study.
@@ 3.2 Collaborative Bandits
pal2024improving address the challenge of optimizing time slot selection in mobile health programs like Kilkari, where the goal is to deliver automated voice messages to beneficiaries at times they are most likely to engage. To tackle this, pal2024improving formulate the problem as a multi-agent multi-armed bandit problem slivkins2019introduction. Here, each beneficiary is modeled as an agent, and each possible time slot for delivering the message is considered an arm. The key idea is to learn the preferences of these agents (beneficiaries) for different arms (time slots) through repeated interactions (i.e., attempting to deliver messages).
To efficiently solve this multi-agent bandit problem, the underlying algorithm in pal2023optimalalgorithmslatentbandits; pal2023optimal is applied to the current problem. This framework leverages the assumption that the preferences of beneficiaries are not entirely independent but rather share some underlying structure. Specifically, it assumes that the matrix representing beneficiaries’ preferences (e.g., the probability of a beneficiary picking up a call at a given time slot) is approximately low-rank. This low-rank assumption implies that there are a few latent factors that explain a significant portion of the variability in beneficiaries’ time slot preferences. The collaborative bandit algorithm exploits this low-rank structure to learn more efficiently by sharing information across beneficiaries, rather than learning each beneficiary’s preferences in isolation.
pal2024improving introduces two novel algorithms: Greedy Matrix Completion (MC) and Phased MC. Phased MC is the key algorithm we use in this work. Greedy MC first has a long random phase where arms are picked randomly, followed by prediction which is followed thereafter. Phased MC operates by updating the estimates in ”phases” which implies that no lengthy exploration phase is required to obtain an initial estimate. To allow for exploration during prediction, a Boltzmann noise cesa2017boltzmann is added. It also uses variance reduction techniques to improve robustness to noise. Since the exploration phase required in this algorithm is not as large, it hence prevents chances of dropout early in the trial due to random calls.
@@ 4.1 Randomization
Random Group (Control): This group received calls using the current Kilkari call retry algorithm Bashingwa2021. This group considered $6416$ beneficiaries for the analysis.
Treatment Group (Collaborative Bandit): This group received calls scheduled using a collaborative bandit algorithm designed to learn and adapt to individual beneficiaries’ preferred call times. This group considered $6490$ beneficiaries for the analysis. The algorithm from pal2024improving relies on an offline low rank matrix completion oracle $\mathcal{O}$. For an unknown matrix $Z\in\mathbb{R}^{m\times n}$, $\mathcal{O}$ takes a subset of noisy observations of a matrix as position $\omega$ ($\{M_{ij}\}_{(i,j)\in\omega}$) as input, and returns an estimate $\hat{Z}$ of $Z$. To implement this oracle, we minimize the following nuclear norm regularized objective:
[…]
where $\lambda>0$ is the regularization parameter and $||\cdot||_{\star}$ denotes the nuclear norm for a matrix as shown in pal2024improving. The hyper-parameters of the model were tuned by holding out 20% of the most recent call data as validation, and then choosing the best hyper-parameters by calculating the score on this data based on predictions via grid search.
@@ 5 Preliminaries
In this section we mathematically define the call pick-up rates and its variants that are used for the analysis in the later section. The index $i$ represents a beneficiary (or user), $j\in[1,7]$ being one of the seven time slot IDs chosen, $t$ being the day, and $r\in[0,2]$ being the re-attempt number for that slot, i.e. $r=0$ being the first call, and $r=1$ being the second call made, if the first call wasn’t picked up. The time windows for the 7 time slot IDs are detailed in the Appendix in Table 6. Let callcall be mapped uniquely to the tuple $(i,j,t,r)$. Let $A_{call}\equiv A_{i,j,t,r}\in\{0,1\}$ denote whether a call attempt was made for user $i$ during time-slot $j$ on day $t$ and whether it was the $r$-th re-attempt. Similarly, let $p_{call}\equiv p_{i,j,t,r}\in\{0,1\}$ denote whether an attempted call was picked or not. We assume that the set of calls $\{call|p_{call}=1\}\subseteq\{call|A_{call}=1\}$. We now define the pooled pick-up rate as,
@@ 6.6 Summary
Some of the generalizable insights learned from this work are as follows. Collaborative bandit algorithms offer a promising approach for personalized intervention delivery in mobile health By iteratively learning from user responses and interactions, these algorithms can adapt to individual preferences and maximize engagement as compared to random calling algorithms. Our goal was to evaluate the algorithm’s ability to improve call pick-up rates compared to the baseline random calling strategy. This research contributes to the growing body of literature in the application of machine learning in mobile health interventions verma2023deployed; mate2022field; nair2022adviser. By demonstrating the effectiveness of a collaborative bandit algorithm in a real-world setting, we highlight the potential for personalized call scheduling to enhance the reach and impact of maternal health programs at scale. Given the national scope of Kilkari and the potential for improved message delivery to millions of beneficiaries, our findings have significant implications for public health policy and practice in India and beyond.